# Quantum Circuit Design using a Progressive Widening Enhanced Monte Carlo Tree Search

### v1

@@ 5 Discussion
The experiments described in previous sections demonstrate that PWMCTS successfully identifies near-optimal ansatz for variational quantum algorithms across various problems. PWMCTS shows a high level of automation given by the robustness of its results across different application domains with the fixed hyperparameter configuration. This robustness is mainly due to the reformulation of the problem and to the progressive widening technique. An adaptive branching factor is important as its optimal value differs substantially within the same domain and depends on the hyperparameters of PWMCTS. On the ground-state energy problem, the molecule of lithium hydride would require a branching factor of $5$ and the molecule of water a branching factor of $50$. Figure 10 shows the dependence of the best cost value on the branching factor. On the left, in blue, is the ground-state energy problem for the Lithium Hydride molecule. On the right, in red, is the ground-state energy problem for the Water molecule. Statistics have been collected over $10$ independent runs for each branching factor value and problem. The scatter plot displays the mean cost values, and the shaded areas indicate the corresponding standard deviations.
@@
Furthermore, we extended the study to an unstructured problem where the goal is to approximate a random quantum state. We experimentally demonstrate that PWMCTS approximates quantum states from a dataset of random quantum circuits. The dataset consists of quantum circuits with $4$, $6$, and $8$ qubits, containing a total number of gates ranging between $5$ and $30$, selected from the set of Clifford generator gates and the CNOT gate. It is important to note that the universal gate set used to generate the dataset of benchmark circuits is different from the one used by PWMCTS. This provides a fair benchmark, free from underlying structural patterns between the benchmark circuits and the search technique. Additionally, we ranked the quantum states generated based on their stabilizer Rényi entropy to analyze the dependence of PWMCTS’s performance on the nonstabilizerness of the target state. PWMCTS manages to design PQCs to approximate quantum states with different values of stabilizer Rényi entropy. The experiments demonstrate that PWMCTS does not exhibit any evident struggles in working with quantum states with high values of nonstabilizerness, supporting PWMCTS as a promising candidate for solving QAS problems and advancing toward quantum advantage. Note that, all the experiments are based on noiseless simulations.
@@ 5.1 Runtime Analysis
The performance analysis of QAS techniques takes into account the value of the cost function on the quantum circuit designed, the count of CNOT gates and parameterized gates, the robustness of the hyperparameter configurations, and runtime. The value of the cost function for the designed PQC measures the quality of the solution to the original problem compared to the optimal cost value. The count of CNOT and parameterized gates reflects the efficiency and complexity of the ansatz. A lower count of CNOT gates is desirable to execute the circuit on NISQ devices reliably, as the error rate of single-qubit gates is typically negligible compared to those of two-qubit gates. The count of parameterized gates affects the optimization difficulty of the ansatz. The robustness of the hyperparameter configuration of QAS techniques measures the degree of automation and adaptability to different problems and domains of application. The runtime analysis is performed in terms of the number of total quantum circuit evaluations $N_{eval}$. It refers to the number of quantum circuits executed on a quantum device to evaluate the cost function until a circuit providing a sufficiently good solution is successfully designed. The number of quantum circuit evaluations required by PWMCTS is
@@ 5.2 Comparison with Nested Monte Carlo Tree Search
We benchmark PWMCTS with the Nested Monte Carlo Tree Search (NMCTS) approach proposed in [18] for two main reasons. First, both techniques inherently represent the QAS problem as a tree and solve it using an MCTS approach with an identical reward function. Second, NMCTS has been extensively tested in various domains and with up to $10$ qubits, making it an appropriate testbed. In contrast, the related works introduced in Section 2 based on different techniques have been primarily tested on small-scale problems within a limited set of application domains.
@@
Table 2 summarizes the experimental results of the two techniques on four problems. The runtime analysis highlights a significant improvement in the number of circuit evaluations required by PWMCTS, achieving a reduction of $10$ to $100$ times compared to NMCTS while delivering the same or better results. The PQCs designed by PWMCTS are also shallower in terms of both CNOT and angle parameter count, up to a factor six and three, respectively. Moreover, PWMCTS achieves these results using a single hyperparameter configuration, while NMCTS requires tuning several hyperparameters for which the values are different for each application.
@@ 6 Conclusions
This article investigated the role of MCTS in designing PQCs for the quantum architecture search problem. The main goal was the development of a problem-agnostic MCTS technique that could reduce the computational cost compared to previous MCTS research [18]. For this purpose, we have proposed PWMCTS, which is a problem-agnostic, hardware-tailored, and gradient-free technique for QAS. It consists of two contributions. The first is a different formulation of the QAS problem in the tree, where each tree node represents a quantum circuit and each edge represents an action that modifies the quantum circuit. These actions are defined by sampling in a set of PQC modifications, which consists of variation in the circuit structure or of a parameter value [19]. The second contribution equips MCTS with a progressive widening technique [20]. It is employed to deal with the infinite number of possible actions, as it adapts the branching factor along the tree. Progressive widening provides MCTS with a higher level of automation reducing the number of hyperparameters to tune across different problems.
@@
The generality of this approach comes not only from the few hyperparameters to be tuned but also from its adaptability to different problems. In principle, it can be employed not only to build up quantum circuits but also for quantum circuit optimization tasks. In fact, by defining an appropriate reward function, we could start the search from the quantum circuit to optimize (the root of the tree) and set the probability distribution that defines the action space such that PWMCTS only searches for shallower circuits.
We identify three main directions for future research. The first is extending PWMCTS to other domains, such as combinatorial optimization and machine learning. For the former, MCTS can be used to design ansatz to solve NP-hard problems, like MaxCut, with a VQE approach. Moreover, it can be employed as a gradient-free technique to optimize the parameters of the PQCs to those problems with a Quantum Approximate Optimization Algorithm [60]. For the latter, PWMCTS might be used to design quantum kernels and also to address overfitting issues in quantum neural networks [61, 62]. The second future direction regards the design and integration of problem-dependent heuristics or machine learning models in the roll-out and expansion steps [48]. In principle, PWMCTS may be applied to any QAS problem. However, as stated by the free-lunch theorem, it cannot compete with independent problem-tailored approaches. It is an open question whether our approach can achieve state-of-the-art results in specific domains by providing heuristics or training machine learning models to guide the search towards more promising states. The third future direction regards MCTS enhancements [43]. Although MCTS enhancements have been extensively studied in the domain of games, their role in the domain of quantum computing is still under investigated. Moreover, quantum enhancements for MCTS are worthy of investigation [63].
@@
The code and supplementary experiments are accessible via Github at the following public repository: https://github.com/VincenzoLipardi/MCTS-QAS.

### v2

@@ 5 Discussion
The experiments described in previous sections demonstrate that PWMCTS successfully identifies near-optimal ansatz for variational quantum algorithms across various problems. PWMCTS shows a high level of automation given by the robustness of its results across different application domains with the fixed hyperparameter configuration. This robustness is mainly due to the reformulation of the problem and to the progressive widening technique.
@@
An adaptive branching factor is important as its optimal value differs substantially within the same domain. In the ground-state energy problem, the lithium hydride molecule would require a branching factor of $5$ and the water molecule a branching factor of $50$. Figure 10 shows the dependence of the best cost value on the branching factor. On the left, in blue, is the ground-state energy problem for the lithium hydride molecule. On the right, in red, is the ground-state energy problem for the water molecule. Statistics have been collected over $10$ independent runs for each branching factor value and for each problem. The scatter plot displays the mean cost values, and the shaded areas indicate the corresponding standard deviations.
The single configuration of hyperparameters, fixed for all the experiments in this article, has been determined with a grid search around reference values found in the literature [19, 59]. We refer the reader to Subsection 3.3 for further details and to Table 1 for the list of hyperparameters and their respective values. This hyperparameter configuration is not guaranteed to be optimal, but it is good enough to obtain relevant results in all test domains without any further problem-dependent hyperparameter search, in contrast to the previous MCTS work [18].
@@
Furthermore, we extended the study to an unstructured problem where the goal is to approximate a random quantum state. We experimentally demonstrate that PWMCTS approximates quantum states from a dataset of random quantum circuits. The dataset consists of quantum circuits with $4$, $6$, and $8$ qubits, containing a total number of gates ranging between $5$ and $30$, selected from the set of Clifford generator gates and the CNOT gate. It is important to note that the universal gate set used to generate the dataset of benchmark circuits is different from the one used by PWMCTS. This provides a fair benchmark, free from the underlying structural patterns between the benchmark circuits and the search technique. Additionally, we ranked the quantum states generated based on their stabilizer Rényi entropy to analyze the dependence of PWMCTS’s performance on the nonstabilizerness of the target state. PWMCTS manages to design PQCs to approximate quantum states with different values of stabilizer Rényi entropy. The experiments demonstrate that PWMCTS does not exhibit any evident struggles in working with quantum states with high values of nonstabilizerness, supporting PWMCTS as a promising candidate for solving QAS problems and advancing toward quantum advantage.
We remark that QAS is a computationally hard problem, as even the training of parameterized quantum circuits is NP-hard [65]. The search space of QAS grows exponentially with the number of qubits, and the scaling of QAS techniques is a key challenge in the field. Although the scaling of PWMCTS to larger quantum circuits and different domains of application may be a potential limitation, it has two main features that help address the scalability problem. The first lies in the balance between exploitation and exploration, originating from the intrinsic nature of MCTS, which aims to confine the search to the branches expected to be more promising rather than searching in the whole, exponentially large, tree. The second is given by the hyperparameter $d$, which restricts the search space by fixing the maximum depth of the quantum circuits explored by PWMCTS.
Note that all the experiments, described in the previous section, are based on noiseless simulations. However, in Section 5.1, we present a preliminary analysis of some experiments under the influence of quantum noise. Section 5.2 describes the metric used to evaluate and compare the efficiency of PWMCTS. Finally, Section 5.3 compares the results obtained by PWMCTS with the state-of-the-art technique on QAS based on MCTS [18].
5.1 Performance Analysis under Noisy Simulations
For NISQ applications, it is important to test the impact of quantum noise on the performance of PWMCTS. In this section, we extend some of the experiments of Section 4 to various simulated quantum channels. For this purpose, we worked on the ground-state energy problem for the hydrogen molecule and on the oracle approximation problem for random quantum circuits introduced in Subsection 4.3.
PWMCTS has been tested on the ground-state energy estimation problem under bit-flip noise [13], depolarizing noise [13], and their combination. These quantum noise channels were chosen because they model noise associated with readout and gate infidelities, respectively [13]. In our experiments, we fix both noise probabilities to $0.01$. Note that this value is one order of magnitude higher than the typical probabilities used to model readout errors and single-qubit gate infidelities in real quantum hardware, and is comparable to those used for two-qubit gate infidelities. The typical values for the noise probabilities are based on the IBM Ourense quantum device, which, although currently unavailable, enables comparison with previous QAS studies that investigated the influence of quantum noise [26, 38]. PWMCTS has been tested on the oracle approximation problem, on circuits with four and six qubits, under different levels of bit-flip noise. Given the larger size of the problem, we did not consider depolarizing noise in this case due to computational constraints. Instead, we used significantly higher bit-flip noise probabilities: $0.1$ and $0.2$.
Figure 11 compares the noiseless results obtained in Section 4 with the results under simulated noise. Figure 11(a), on the left, plots the energy values for all the noise models tested and for different values of PWMCTS iterations $I$ ($I=1000$ in green, $I=10000$ in blue, $I=100000$ in red). The circles indicate the results of the best node (quantum circuit) in the tree, while the squares represent the results after the fine-tuning of the parameters with the Adam optimizer. Figure 11(a) shows that PWMCTS does not manage to compete with the results of the classical SCF and FCI benchmarks. However, increasing the number of PWMCTS simulations significantly improves the results. As expected, the worst results are obtained on the noise model that combines bit-flip and depolarizing noise. On the one hand, the results obtained by PWMCTS with $I=10{,}000$ and $I=100{,}000$ improve upon those reported in [26]. On the other hand, the reinforcement learning approach proposed for QAS in [38] manages to reach the same energy as FCI within chemical accuracy. However, this latter approach has only been tested in the quantum chemistry domain and incorporates domain knowledge through machine learning. In contrast, our approach is completely domain-agnostic. We note that, for a given problem, the number of gates and parameters required in the quantum circuit to solve it is expected to increase with the noise level [66, 67, 38]. Therefore, in our case, to find the optimal circuit in the presence of noise, we should increase the value of the hyperparameter $d$, which constrains the depth of the circuit.
Figure 11(b), on the right, shows the influence of bit-flip noise on the performance of PWMCTS in approximating four of the quantum circuits defined in Section 4.3. Specifically, we considered the four circuits with the number of gates fixed to $20$, denoted as ($n=\{4,6\}$, $g=20$, $M_{2}=\{\text{easy},\text{hard}\}$). Performance is measured in terms of the number of successful approximations according to the definition given in Section 4.3.1, with the same $\epsilon$ values. The first column in Figure 11(b) reports the same results shown in Figure 9, while the second and third columns refer to experiments with increasing bit-flip noise probability. The first two rows indicate that for the easiest problems, $4$-qubit quantum circuits, PWMCTS is significantly robust. It only fails to approximate the circuit in 1 out of 10 independent run ($4$, $20$, hard) under the highest noise level. Similarly, the last row, related to the 6-qubit circuit with the highest nonstabilizerness, shows that PWMCTS obtains the same number of successful approximation as in the noiseless case and it only decreases of one for the highest noise level.
(a) H2
(b) Random quantum circuits
Figure 11: Comparison of PWMCTS performance under various quantum noise channels. Figure 11(a) presents the results for the ground-state energy estimation of the hydrogen molecule across different values of PWMCTS iterations and noise channels. Figure 11(b) illustrates the performance on the oracle approximation problem for four random quantum circuits, defined in Subsection 4.3, comparing different levels of bit-flip noise.
QAS in the noiseless scenario can be considered a single-player game in deterministic environments. In the presence of noise multiple evaluations of the cost function on the same circuit may bring different values. QAS in the noisy scenario can be considered a single-player game in a stochastic environment. We remark that the results obtained in this section are based on exactly the same implementation of PWMCTS, including the hyperparameter configuration. However, PWMCTS would benefit from MCTS refinements to better cope with this change. Several MCTS variants have been proposed to handle stochastic single-player games [45, 46] and we believe it is worthy to further investigate in this direction.
We emphasize that PWMCTS requires further studies before it can be reliably used to design quantum circuits for real quantum hardware. Nevertheless, two main reasons motivate the consideration of PWMCTS as a promising candidate for tackling QAS on real hardware. First, there is a need of gradient-free techniques, which are usually more robust to noise-induced barren plateaus [68]. Second, extensive research has already been conducted in the classic domain of games, where MCTS behaves well even in stochastic settings [45, 46]. The results presented in this article are promising in this regard, especially considering that we did not incorporate any enhancements to MCTS for handling stochastic environments.
@@ 5.2 Runtime Analysis
The performance analysis of QAS techniques takes into account the value of the cost function on the quantum circuit designed, the count of CNOT gates and parameterized gates, the robustness of the hyperparameter configurations, and runtime. The value of the cost function for the designed PQC measures the quality of the solution to the original problem compared to the optimal cost value. The count of CNOT and parameterized gates reflects the efficiency and complexity of the ansatz. A lower count of CNOT gates is desirable to execute the circuit on NISQ devices reliably, as the error rate of single-qubit gates is typically negligible compared to that of two-qubit gates. The count of parameterized gates affects the optimization difficulty of the ansatz. The robustness of the hyperparameter configuration of QAS techniques measures the degree of automation and adaptability to different problems and domains of application.
The runtime analysis is performed in terms of the number of total quantum circuit evaluations $N_{eval}$. It refers to the number of quantum circuits executed on a quantum device to evaluate the cost function until a circuit providing a sufficiently good solution is successfully designed. The number of quantum circuit evaluations required by PWMCTS is
@@
According to previous work on QAS [18], the runtime analysis is based on the number of quantum circuit evaluations. This serves as the reference comparison metric, as it is independent of hardware, software libraries, and the specific quantum circuits explored during the search. We note that the progressive widening technique [20] does not result in significant computational overhead for MCTS.
@@ 5.3 Comparison with Nested Monte Carlo Tree Search
We benchmark PWMCTS with the Nested Monte Carlo Tree Search (NMCTS) approach proposed in [18] for two main reasons. First, both techniques inherently represent the QAS problem as a tree and solve it using an MCTS approach with an identical reward function. Second, NMCTS has been extensively tested in various domains and with up to $10$ qubits, making it an appropriate testbed. In contrast, the related works introduced in Section 2 based on different techniques have been mainly tested on small-scale problems within a limited set of application domains.
@@
Table 2 summarizes the experimental results of the two techniques on four problems. The runtime analysis highlights a significant improvement in the number of circuit evaluations required by PWMCTS, achieving a reduction of $10$ to $100$ times compared to NMCTS while delivering the same or better results. The PQCs designed by PWMCTS are also shallower in terms of both CNOT and angle parameter count, up to a factor 6 and 3, respectively. Moreover, PWMCTS achieves these results using a single hyperparameter configuration, while NMCTS requires tuning several hyperparameters, for which the values are different for each application.
@@ 6 Conclusions
This article investigated the role of MCTS in designing PQCs for the quantum architecture search problem. The main goal was the development of a problem-agnostic MCTS technique that could reduce the computational cost compared to previous MCTS research [18]. For this purpose, we have proposed PWMCTS, which is a problem-agnostic, hardware-tailored, and gradient-free technique for QAS. It consists of two contributions. The first is a different formulation of the QAS problem in the tree, where each tree node represents a quantum circuit and each edge represents an action that modifies the quantum circuit. These actions are defined by sampling in a set of PQC modifications, which consists of variation in the circuit structure or of a parameter value [19]. The second contribution equips MCTS with a progressive widening technique [20]. It is employed to deal with an infinite number of possible actions, as it adapts the branching factor along the tree. Progressive widening provides MCTS with a higher level of automation reducing the number of hyperparameters to tune across different problems.
@@
The generality of this approach comes not only from the few hyperparameters to be tuned, but also from its adaptability to different problems. In principle, it can be employed not only to build up quantum circuits but also for quantum circuit optimization tasks [69]. In fact, by defining an appropriate reward function, we could start the search from the quantum circuit to optimize (the root of the tree) and set the probability distribution that defines the action space such that PWMCTS only searches for shallower circuits.
We identify four main directions for future research. The first deals with the robustness PWMCTS under the influence of quantum noise. The preliminary experiments show promising results for the performance of PWMCTS under various noise models, both for the ground-state energy estimation of the hydrogen molecule and for the oracle approximation problem. However, it is an interesting direction to include MCTS enhancements developed for stochastic settings [45, 46], to QAS. Moreover, we also highlight the implementation of PWMCTS on real quantum hardware as a future direction. The second future direction is extending PWMCTS to other domains, such as combinatorial optimization and machine learning. For the former, MCTS can be used to design ansatz to solve NP-hard problems, like MaxCut, with a VQE approach. Moreover, it can be used as a gradient-free technique to optimize the parameters of the PQCs to those problems with a Quantum Approximate Optimization Algorithm [70]. For the latter, PWMCTS might be used to design quantum kernels and also to address overfitting issues in quantum neural networks [71, 72]. The third future direction deals with the design and integration of problem-dependent heuristics or machine learning models in the roll-out and expansion steps [51]. In principle, PWMCTS may be applied to any QAS problem. However, as stated by the free-lunch theorem, it cannot compete with independent problem-tailored approaches. It is an open question whether our approach can achieve state-of-the-art results in specific domains by providing heuristics or training machine learning models to guide the search towards more promising states. The fourth future direction deals with the MCTS enhancements [46]. Although MCTS enhancements have been extensively studied in the domain of games, their role in the domain of quantum computing is still not well investigated. Moreover, quantum enhancements for MCTS are worthy of investigation [73].
@@
The code and supplementary experiments are accessible via Github and the source code archived in Zenodo at time of publication.
Acknowledgements
The authors acknowledge the anonymous reviewers for their helpful comments and suggestions that improved the manuscript, especially regarding the experiments with simulated noise.